OpenStudy (anonymous):

Find fraction notation for the ratio 68 to 1.38

myininaya (myininaya):

\[\frac{68}{1.38}*\frac{100}{100}=\frac{6800}{138}\] you can simplify

OpenStudy (anonymous):

how sure are you of the answer

OpenStudy (mattfeury):

well \( \frac{100}{100} = 1\) and anything times 1 is itself, isn't it?

OpenStudy (anonymous):

i am not good on fraction i have some mre can you help me more total is 30

myininaya (myininaya):

what does that mean? more total is 30?

OpenStudy (anonymous):

i have about 30 of them to do

myininaya (myininaya):

the reason i multiply by 100/100 is so i could move that decimal over twice in the bottom

OpenStudy (anonymous):

find fraction notation for the ratio 73 to 100

myininaya (myininaya):

and like matt said 100/100 is 1 so multiplying by 1 does not change the value of the fraction since 1 is the multiplicative identity

myininaya (myininaya):

\[\frac{73}{100}\]

myininaya (myininaya):

if you have find fraction notation for the ratio a to b just write it as a/b